OpenStudy (anonymous):

2x4(4x2 + 3x + 1) multiplying polynomials help please!!!

OpenStudy (anonymous):

hint... \[\LARGE a(b+c+d)=ab+ac+ad\] and also \[\LARGE x^a\cdot x^b=x^{a+b}\] ...

OpenStudy (anonymous):

another hint???

OpenStudy (anonymous):

\[\Large 2x^4(4x^2+3x+1)=2x^4\cdot 4x^2+2x^4\cdot 3x+2x^4\cdot 1 \]

OpenStudy (anonymous):

need more help? :)

OpenStudy (anonymous):

nope thanks can take it form there

OpenStudy (anonymous):

My pleasure, Good Luck.